find an equation of the line containing the given points and having the given slope. Write the equation in slope-intercept
(6,-3), m=-4
(4,2),m-2/3

Answer by Boreal(15235)

point slope formula
y-y1=m(x-x1), m is slope (x1,y1) point
y+3=-4(x-6)
y+3=-4x+24
y=-4x+21
--------------
similarly, y-2=(2/3)(x-4)
y=(2/3)x-(8/3)+6/3
y=(2/3)x-(2/3)
Note, on this one I am assuming m=2/3. if m is -2/3, then the sign has to change and the result will be
y=(-2/3)x+(14/3)